OpenStudy (anonymous):

What is the slope of a line that passes through the point (−2, 3) and is parallel to a line that passes through (3, 7) and (−2, −8)? (4 points)

OpenStudy (binary3i):

y = 3x + 9

OpenStudy (anonymous):

find the slope between (3,7) and (-2, -7) using (y1-y2)/(x1-x2) = m use that slope and plug it into point-slope form: y-y1=m(x-x1) where x1,y1 are the coordinates of the given point.

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

can you calculate the slope between (3,7) and (-2,-8) as i described?

OpenStudy (anonymous):

m = slope

OpenStudy (anonymous):

i still dont' understand why m stands for slope? was slope named something different before?

OpenStudy (anonymous):

mlope?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

omg i just dont understand at all

OpenStudy (anonymous):

okay... see these two point: (3,7) and (-2,-8) they have an x coordinate (the first number) and a y coordinate (the second number) subtract the second number from each point and divide the difference of the x coordinates.

OpenStudy (anonymous):

second numbers (y coordinates) -8, 7. subtract them: -8 - 7 = -15 first numbers (x coordinates): -2, 3. subtract them: -2 - 3 = -5 divide them: -15/-5 = 3 that is your slope.

OpenStudy (anonymous):

any help?

OpenStudy (anonymous):

yes thank you :)

OpenStudy (anonymous):

ok.. with that, you need to put it in poin-slope form: y - y1 = m(x - x1) you already have your slope m = 3. the given point is (-2, 3). just plug those values in to the formula i just stated.